Tohban Report 2018-04-04
From RHESSI Wiki
|Start Date:||28 Mar 2018|
|End Date:||4 Apr 2018|
|Tohban:||Juan Carlos Martinez Oliveros|
|Next Tohban:||Mitsuo Oka|
|List all reports|
Solar activity this week was relatively low. There are not visible active regions on disk. Solar activity is expected to be very low for the next three days (04-06 Apr).
TIME RANGE:28-Mar-18 00:00:00.000 -- 04-Apr-18 00:00:00.000 % RD_WEEK_FILE: No input file for week 18_13 found in: $DIR_GEN_GEV % RD_WEEK_FILE: No input file for week 18_14 found in: $DIR_GEN_GEV No GOES flares for Time Range % HSI_QLOOK::ONLY_TIME_RANGE: No Flares found % HSI_TOHBAN_FLARE_COUNTER: No RHESSI Flare list for this time period % HSI_TOHBAN_FLARE_COUNTER: No RHESSI Flares for this time period
The SSR reached a maximum of ~95 percent during this week.
Detectors 3 and 8 are active and taking data.
The cold plate 1 and 2 temperatures are ~156.4K and 155.3K. As last week.
Currently the Cryocooler power is ~55 W and the acceleration is at ~11mg
= Detector issues
6 hours of VC1 data due to poor communications at AGO.
Data was deleted on March 31 as the SSR fill level was very high. This move brought the SSR fill level down from 95% to 65% full. We skipped over data recorded from 2018-082-08:36 to 2018-084-12:07 UTC.
A spin down experiment is currently been make. The expectation is that reducing the mechanical stress on the s-band communication subsystem, stable communications with the spacecraft can be achieved.
The procedure started on 3 April 2018 at 23:00 UT. The spacecraft achieved 10 RPM on 4 Apr 2018.
|HLAT Decimation||Rear decimation|
|Night time data (fronts)||plus/minus 4 minutes from daylight|
|Night time data (rears)||plus/minus 4 minutes from daylight|
|Require extra passes?||No|
|Requirement for moving pointer?||Yes|
|Attenuator operation||logic enabled|
|Detector problems?||See notes above.|